# Stereo4D: Learning How Things Move in 3D from Internet Stereo Videos

### v1

@@ Abstract
Learning to understand dynamic 3D scenes from imagery is crucial for applications ranging from robotics to scene reconstruction. Yet, unlike other problems where large-scale supervised training has enabled rapid progress, directly supervising methods for recovering 3D motion remains challenging due to the fundamental difficulty of obtaining ground truth annotations. We present a system for mining high-quality 4D reconstructions from internet stereoscopic, wide-angle videos. Our system fuses and filters the outputs of camera pose estimation, stereo depth estimation, and temporal tracking methods into high-quality dynamic 3D reconstructions. We use this method to generate large-scale data in the form of world-consistent, pseudo-metric 3D point clouds with long-term motion trajectories. We demonstrate the utility of this data by training a variant of DUSt3R to predict structure and 3D motion from real-world image pairs, showing that training on our reconstructed data enables generalization to diverse real-world scenes. Project page: https://stereo4d.github.io
@@ 1 Introduction
Simultaneously predicting and understanding geometry and motion—that is, dynamic 3D content—from images is a fundamental building block for computer vision, with applications ranging from robotic interaction and scene reconstruction to novel view synthesis of dynamic scenes. While recent work has made remarkable progress in predicting static 3D structure from images [99, 105, 5], modeling real-world 3D motion—people gesturing, balls bouncing, leaves rustling in the wind—remains a critical unsolved challenge for building truly general models of the visual world.
Recent breakthroughs in AI, from large language models [1, 89] to image generation [73] and static 3D reconstruction [99, 5, 105], demonstrate a consistent pattern: large amounts of high-quality, realistic training data and scalable architectures enable dramatic performance improvements. In the realm of 3D reasoning, prior works [49, 104, 74, 75, 99] have shown the value of large-scale training data for strong zero-shot generalization within single-view or two-view static scene settings. But applying this same formula to dynamic 3D scenes (i.e. moving 3D structure) requires a corresponding large-scale dataset consisting of diverse visual content paired with corresponding ground-truth 3D motion trajectories. Obtaining such data presents unique challenges. While there are synthetic datasets [115, 9, 19, 29], these often fail to capture the distribution of real-world content and the nuanced patterns of real-world motion. Traditional approaches to gathering real motion data, such as motion capture systems or multi-view camera arrays [38, 28, 43, 35] are accurate, but difficult to scale and limited in the diversity of scenes they can capture.
@@
Our contributions include: (1) a framework for obtaining real-world, dynamic, and pseudo-metric 4D reconstructions and camera poses at scale from existing online video; (2) DynaDUSt3R, a method that takes a pair of frames from any real-world video, and predicts a pair of 3D point clouds and the corresponding 3D motion trajectories that connect them in time.
@@ 2 Related work
2D and 3D motion data. There has been tremendous progress on the task of motion estimation from images and videos, and in particular for 2D image-space correspondence estimation. Most state-of-the-art methods use neural networks trained on ground truth data to predict these correspondences directly from images. While these approaches require large training datasets, synthetic data from graphics engines [19, 59, 30, 9, 83, 29, 115] has proven surprisingly effective at generalizing to real-world data, likely because the core task, low-level textural correspondence, is similar between the two domains.
@@
However, the same cannot be said for 3D motion estimation, since predicting both 3D structure and motion is usually more ambiguous and can require specific prior knowledge about the real world and how it moves. To help address this domain gap, a number of real-world datasets have been proposed. The KITTI [27] and Waymo [86] datasets include real-world autonomous driving sequences with stereo and motion annotations derived from LiDAR and odometry information, but only focus on the relatively closed domain of street scenes, whereas our data depicts more diverse in-the-wild scenarios. A number of annotated smaller-scale datasets, such as TAPVid [16], TAPVid3D [46], and Dycheck [25], have been proposed, primarily serving as evaluation datasets for benchmarking depth estimation, 3D reconstruction, and 3D motion estimation approaches. WSVD [97] and NVDS [100] are stereo video datasets that include disparity maps derived from optical flow. While their source content is similar, our method provides richer 3D annotations beyond time-independent disparity maps, such as 3D camera parameters and long-term 3D motion tracks.
@@
Static and dynamic scene reconstruction The problem of reconstructing a static 3D scene has been studied for decades. Traditional 3D reconstruction methods tackle this problem by first estimating camera parameters via Structure-from-Motion (SfM) [82, 72, 71, 72, 2, 76, 87, 32] or SLAM [20, 11, 60, 15]. Dense scene geometry can then be estimated through Multi-view Stereo (MVS) [10, 36, 23, 22, 24, 77, 106, 107] followed by surface reconstruction algorithms [33, 14, 40]. More recently, deep neural network-based approaches have shown promising results in improving camera localization accuracy or scene reconstruction through intermediate representations such as depth maps [4, 88, 93, 91, 79, 54], radiance fields [55, 21, 69, 26, 80, 102], or 3D scene coordinates [7, 8, 48, 99, 111]. However, these methods assume the input images to be observations of a static environment, and therefore produce inaccurate geometry and camera poses for dynamic scenes.
@@
Reconstructing dynamic scenes is more challenging since scene and object motions violate the multi-view constraints used to reconstruct static scenes. As a result, many prior works require RGBD input [6, 62] or only recover sparse geometry [66, 96, 81]. Several recent works tackle this problem from monocular input through video depth maps [113, 45, 114], time-varying radiance fields [67, 68, 51, 56, 52, 25, 47, 98], or generative priors [103].
@@ 3 Creating a dataset of 4D scenes
A core contribution of this work is a pipeline for extracting high-quality, pseudo-metric, 3D data from online stereoscopic fisheye videos (known as VR180 videos). High-resolution, wide field of view VR180 videos can be found readily online. We show that this data is ideal for deriving rich dynamic 3D information that can power models for predicting geometry and motion from imagery.
@@ 3.1 Data Processing Pipeline
At a high level, our pipeline for converting a stereoscopic video into a dynamic point cloud involves estimating camera poses, stereo disparity, and 2D tracks, fusing these quantities into a consistent 3D coordinate frame, and performing several filtering operations to ensure temporal consistent, high-quality reconstructions (Fig. 2). In this section, we describe in detail the key components of this process.
@@
Depth Estimation. We next estimate a per-frame disparity map, operating on each frame independently. In particular, we use the estimated camera rig calibration $\mathbf{c}_{r},\mathbf{R}_{r}$ to create rectified stereo pairs from the stereo fisheye video and estimate the per-frame disparity $\mathbf{D}_{i}$ with RAFT [84, 83, 90].
3D Track Estimation and Optimization. We extract long-range 2D point trajectories using BootsTAP [17]. Using the camera poses $\mathbf{c}_{i},\mathbf{R}_{i}$ and disparity maps $\mathbf{D}_{i}$, we unproject these tracked points into 3D space, turning each 2D track $j$ into a 3D motion trajectory $\mathbf{p}^{j}_{1},\ldots,\mathbf{p}^{j}_{N}$ across all frames. In general, each point will usually only be tracked in a subset of frames, but for simplicity, we describe the formulation as if all points are always visible. Moreover, since subsequent steps are done independently per-track, we drop the superscript $j$ in future references.
@@
Figure 3: Effect of track optimization. Comparing motion trajectories before and after track optimization, we see that optimization resolves the high frequency jitter along camera rays, affecting both static and dynamic content. After optimization, static content has static tracks, and dynamic tracks are less noisy.
Figure 4: Diverse motion: Stereo4D captures a wide variety of types of moving objects, from swimming fish, to walking pedestrians, moving vehicles, and a farmer sowing seeds. It includes source videos captured with both stationary (left) and moving (right) cameras.
Implementation details. Shot-selection. Rather than work with the full video, we break the footage into discrete, trackable shots using ORB-SLAM2’s stereo estimation mode [61] following [118]. Field of View. While estimating pose, we use a $140^{\circ}$ FoV fisheye format, which we found to capture more of the (usually static) background and less of the (often dynamic) foreground, yielding more reliable camera poses. Stereo Confidence Checks. We discard pixels where the $y$-component of RAFT flow is more than 1 pixel (since rectified stereo pairs should have perfectly horizontal motion) and where the stereo cycle consistency error is more than 1 pixel (since such pixels are unreliable). Dense 2D tracks. To get dense tracks, we run BootsTAP with dense query points: for every 10th frame, we uniformly initialize $128\times 128$ query points on frames of resolution 512 $\times$ 512. We then prune redundant tracks that overlap on the same pixel. Drifting tracks. Since 2D tracks can drift on textureless regions, we discard moving 3D tracks that correspond to certain semantic categories (e.g., “walls”, “building”, “road”, “earth”, “sidewalk”), detected by DeepLabv3 [13] on ADE20K classes [116, 117].
Filtering details. A fraction of the video clips that are processed may be unsuitable because they either (1) are not videos, and are entirely static images, (2) contain cross-fades, or (3) have text or other synthetic graphics. To discard text and title sequences, we avoid creating video clips from the start and ends of the source videos. We identify cross-fades by running SIFT [57] matching through the video at multiple temporal scales and discarding video clips with static camera but with fewer than 5 SIFT matches between frames that are 5 seconds apart.
@@ 3.2 Stereo4D Dataset
Fig. 4 illustrates a subset of videos and reconstructions from a dataset processed with the above pipeline, encompassing more than 100K clips capturing everyday scenes and activities. To visualize the range of content, we used an automatic captioning system to generate captions for the dataset and created a word cloud (Fig. 5) highlighting the most frequently observed objects.
@@ 5 Experiments
We conduct a series of experiments to validate the effectiveness of our proposed data and techniques. First, we evaluate our proposed real-world Stereo4D data mined from VR180 videos on the DynaDUSt3R task. In particular, we compare models that are individually trained with our real-world data and with synthetic data, and we show that our data enables model learning more accurate 3D motion priors (Sec. 5.1). Second, we show that our trained model that adapts DUSt3R has strong generalization to in-the-wild images of dynamic scenes, and enables accurate predictions of underlying geometry (Sec. 5.2).
@@ 7 Stereo4D Statistics
We collected around 110k clips from 6,493 Internet VR180 videos. Fig. 12 shows the camera translation distribution between the first and last frame of each clip. In Fig. 12, we measure the motion in terms of pixel displacement projected onto the image frame. Measuring motion in pixel-space emphasizes motion that occurs closer to the camera, since such motion yields larger pixel displacements, while naturally de-emphasizing motion further from the camera.
@@ 8.1 More results on held-out Stereo4D examples
Fig. 13 shows additional DynaDUSt3R predictions on the Stereo4D held-out test set, extending Fig. 7 from the main paper. Fig. 14 shows additional qualitative examples of motion comparisons on Stereo4D test set, extending Fig. 8 from the main paper. Fig. 14 compares variants of DynaDUSt3R trained on different data sources. The model trained on PointOdyssey incorrectly predicts large 3D motions, while the model trained on Stereo4D makes more accurate motion predictions, closer to ground truth.
Figure 12: Scene motion statistics from Stereo4D. We measure the motion in terms of pixel displacement projected onto the image frame. For each video, we measure the percentage of tracks that have 3D trail length greater than 50 pixels. The 3D trail length is measured by Eqn. 3.
8.2 More qualitative examples on track optimization

### v2

@@ Abstract
Learning to understand dynamic 3D scenes from imagery is crucial for applications ranging from robotics to scene reconstruction. Yet, unlike other problems where large-scale supervised training has enabled rapid progress, directly supervising methods for recovering 3D motion remains challenging due to the fundamental difficulty of obtaining ground truth annotations. We present a system for mining high-quality 4D reconstructions from internet stereoscopic, wide-angle videos. Our system fuses and filters the outputs of camera pose estimation, stereo depth estimation, and temporal tracking methods into high-quality dynamic 3D reconstructions. We use this method to generate large-scale data in the form of world-consistent, pseudo-metric 3D point clouds with long-term motion trajectories. We demonstrate the utility of this data by training a variant of DUSt3R to predict structure and 3D motion from real-world image pairs, showing that training on our reconstructed data enables generalization to diverse real-world scenes. Project page and data at https://stereo4d.github.io
@@ 1 Introduction
Simultaneously predicting and understanding geometry and motion—that is, dynamic 3D content—from images is a fundamental building block for computer vision, with applications ranging from robotic interaction and scene reconstruction to novel view synthesis of dynamic scenes. Recent work has made remarkable progress in predicting static 3D structure from images [100, 106, 5], but modeling real-world 3D motion—people gesturing, balls bouncing, leaves rustling in the wind—remains a critical unsolved challenge for building truly general models of the visual world.
Recent breakthroughs in AI, from large language models [1, 89] to image generation [73] and static 3D reconstruction [100, 5, 106], demonstrate a consistent pattern: large amounts of high-quality, realistic training data and scalable architectures enable dramatic performance improvements. In the realm of 3D reasoning, prior works [49, 105, 74, 75, 100] have shown the value of large-scale training data for strong zero-shot generalization within single-view or two-view static scene settings. But applying this same formula to dynamic 3D scenes (i.e., moving 3D structure) requires a corresponding large-scale dataset consisting of diverse visual content paired with corresponding ground-truth 3D motion trajectories. Obtaining such data presents unique challenges. While there are synthetic datasets [116, 9, 19, 29], these often fail to capture the distribution of real-world content and the nuanced patterns of real-world motion. Traditional approaches to gathering real motion data, such as motion capture systems or multi-view camera arrays [38, 28, 43, 35] are accurate, but difficult to scale and limited in the diversity of scenes they can capture.
@@
Our contributions include: (1) a framework to obtain real-world, dynamic, and pseudo-metric 4D reconstructions and camera poses at scale from online videos; and (2) DynaDUSt3R, a method that, given a pair of video frames, predicts a pair of 3D point clouds and the corresponding 3D motion trajectories that connect them in time.
@@ 2 Related work
2D and 3D motion data. There has been tremendous progress on motion estimation from images and videos, and in particular for 2D image-space correspondence estimation. Most state-of-the-art methods use neural networks trained on ground truth data to predict these correspondences directly from images. While these approaches require large training datasets, synthetic data from graphics engines [19, 60, 30, 9, 83, 29, 116] has proven surprisingly effective at generalizing to real-world data, likely because the core task, low-level textural correspondence, is similar between the two domains.
@@
However, the same cannot be said for 3D motion estimation, since predicting both 3D geometry and motion is often more ambiguous and can require specific prior knowledge about the real world and how it moves. To help address this domain gap, a number of real-world datasets have been proposed. The KITTI [27] and Waymo [86] datasets include real-world autonomous driving sequences with stereo and motion annotations derived from LiDAR and odometry information, but focus on the relatively closed domain of street scenes. Our data depicts more diverse in-the-wild scenarios. A number of annotated smaller-scale datasets, such as TAPVid [16], TAPVid3D [46], and Dycheck [25], have been proposed, primarily serving as evaluation datasets for benchmarking depth estimation, 3D reconstruction, and 3D motion estimation approaches. WSVD [97] and NVDS [101] are stereo video datasets that include disparity maps derived from optical flow. While their source content is similar, our method provides richer 3D annotations beyond time-independent disparity maps, such as camera parameters and long-term 3D motion tracks.
@@
Static and dynamic scene reconstruction The problem of reconstructing a static 3D scene has been studied for decades. Traditional 3D reconstruction methods tackle this problem by first estimating camera parameters via Structure-from-Motion (SfM) [82, 72, 71, 2, 76, 87, 32] or SLAM [20, 11, 61, 15]. Dense scene geometry can then be estimated through Multi-view Stereo (MVS) [10, 36, 23, 22, 24, 77, 107, 108] followed by surface reconstruction algorithms [33, 14, 40]. More recently, deep neural network-based approaches have shown promising results in improving camera localization accuracy or scene reconstruction through intermediate representations such as depth maps [4, 88, 93, 91, 79, 54], point tracks [98, 56], radiance fields [55, 21, 69, 26, 80, 103], or 3D scene coordinates [7, 8, 48, 100, 112]. However, these methods assume the input images to be observations of a static environment, and therefore produce inaccurate geometry and camera poses for dynamic scenes.
@@
Reconstructing dynamic scenes is more challenging since scene and object motions violate the multi-view constraints used to reconstruct static scenes. As a result, many prior works require RGBD input [6, 62] or only recover sparse geometry [66, 96, 81]. Recent works tackle this problem from monocular input through video depth maps [114, 45, 115], time-varying radiance fields [67, 68, 51, 57, 52, 25, 47, 99], or generative priors [104].
@@ 3 Creating a dataset of 4D scenes
A core contribution of our work is a pipeline for extracting high-quality, pseudo-metric, 3D data from online stereoscopic fisheye videos (known as VR180 videos). High-resolution, wide field of view VR180 videos can be found readily online. We show that this data is ideal for deriving rich dynamic 3D information that can power models for predicting geometry and motion from imagery.
@@ 3.1 Data Processing Pipeline
At a high level, our pipeline for converting a stereoscopic video into a dynamic point cloud involves estimating camera poses, stereo disparity, and 2D tracks, fusing these quantities into a consistent 3D coordinate frame, and performing several filtering operations to ensure temporal consistent, high-quality reconstructions (Fig. 2). In this section, we describe the key components of this process.
@@
Depth Estimation. We next estimate a per-frame disparity map, operating on each frame independently. In particular, we use the estimated camera rig calibration $\mathbf{c}_{r},\mathbf{R}_{r}$ to create rectified stereo pairs from the stereo fisheye video and estimate per-frame disparity $\mathbf{D}_{i}$ with RAFT [84, 83, 90].
3D Track Estimation and Optimization. We extract long-range 2D point trajectories with BootsTAP [17]. Using the camera poses $\mathbf{c}_{i},\mathbf{R}_{i}$ and disparity maps $\mathbf{D}_{i}$, we unproject these tracked points into 3D space, turning each 2D track $j$ into a 3D motion trajectory $\mathbf{p}^{j}_{1},\ldots,\mathbf{p}^{j}_{N}$ across all frames. In general, each point will usually only be tracked in a subset of frames, but for simplicity, we describe the formulation as if all points are always visible. Moreover, since subsequent steps are done independently per-track, we drop the superscript $j$ in future references.
@@
Figure 3: Effect of track optimization. Comparing motion trajectories before and after track optimization, we see that optimization resolves the high-frequency jitter along camera rays, affecting both static and dynamic content. After optimization, static content has static tracks, and dynamic tracks are less noisy.
Figure 4: Diverse motion: Stereo4D captures a wide variety of types of moving objects: swimming fish, walking pedestrians, moving vehicles, a farmer sowing seeds, etc. It includes source videos captured with both stationary (left) and moving (right) cameras.
Implementation details. Shot-selection. Rather than work with the full video, we break the footage into discrete, trackable shots using ORB-SLAM2’s stereo estimation mode [61] following [119]. Field of view. While estimating pose, we use a $140^{\circ}$ FoV fisheye format, which we found to capture more of the (usually static) background and less of the (often dynamic) foreground, yielding more reliable camera poses. Stereo confidence checks. We discard pixels where the $y$-component of RAFT flow is more than 1 pixel (since rectified stereo pairs should have perfectly horizontal motion) and where the stereo cycle consistency error is more than 1 pixel (since such pixels are unreliable). Dense 2D tracks. To obtain dense tracks, we run BootsTAP with dense query points: for every 10th frame, we uniformly initialize $128\times 128$ query points on frames of resolution 512 $\times$ 512. We then prune redundant tracks that overlap on the same pixel. Drifting tracks. Since 2D tracks can drift on textureless regions, we discard moving 3D tracks that correspond to certain semantic categories (e.g., “walls”, “building”, “road”, “earth”, “sidewalk”), detected by DeepLabv3 [13] on ADE20K classes [117, 118].
Filtering details. A fraction of the video clips that are processed are unsuitable because they either (1) are not videos, but entirely static images, (2) contain cross-fades, or (3) have text or other synthetic graphics. To discard text and title sequences, we avoid creating video clips from the start and ends of the source videos. We identify cross-fades by running SIFT [58] matching through the video at multiple temporal scales and discarding video clips with static camera but with fewer than 5 SIFT matches between frames that are 5 seconds apart.
@@ 3.2 Stereo4D Dataset
Fig. 4 illustrates a subset of videos and reconstructions from a dataset processed with the above pipeline, encompassing more than 100K clips capturing everyday scenes and activities. To visualize the range of content, we used an automatic captioning system to generate captions for the dataset and created a word cloud (Fig. 5) highlighting the most frequently observed objects. See supplementary for more statistics and ablations on track optimization.
@@ 5 Experiments
We conduct a series of experiments to validate the effectiveness of our proposed data and techniques. First, we evaluate our proposed real-world Stereo4D data mined from VR180 videos on the DynaDUSt3R task. In particular, we compare models that are individually trained with our real-world data and with synthetic data, and we show that our data enables model learning more accurate 3D motion priors (Sec. 5.1). Second, we show that our trained model that adapts DUSt3R has strong generalization to in-the-wild images of dynamic scenes, and enables accurate predictions of underlying geometry (Sec. 5.2). See supplementary for more ablations on DynaDUSt3R design.
@@ 7 Stereo4D Statistics
We collected around 110k clips from 6,493 Internet VR180 videos. The videos were curated from public YouTube content using the tag “VR180”, available under YouTube’s Standard License. We have released derived geometry, motion data, and video links, under a CC license here.
Fig. 13 shows the camera translation distribution between the first and last frame of each clip. Fig. 13 shows, from left to right: (a) the distribution of $x$,$y$,$z$ camera translations (log scale), (b) the distribution of rotations, (c) a sample of 5000 camera trajectories, viewed from above, colored by final camera orientation (red=right, cyan=left).
In Fig. 13, we measure the motion in terms of pixel displacement projected onto the image frame. Measuring motion in pixel-space emphasizes motion that occurs closer to the camera, since such motion yields larger pixel displacements, while naturally de-emphasizing motion further from the camera.
Figure 12: Camera statistics from Stereo4D. From left to right: a) the distribution of x,y,z camera translations (log scale), (b) the distribution of rotations, (c) a sample of 5000 camera trajectories, viewed from above, colored by final camera orientation (red=right, cyan=left)
Figure 13: Scene motion statistics from Stereo4D. We measure scene motion in terms of pixel displacement projected onto the image frame. For each video, we measure the percentage of tracks that have 3D trail length greater than 50 pixels. The 3D trail length is measured by Eqn. 3.
8 Ablations
Figure 14: Model performance with different input time gap: Motion error vs. input frame gap. As the time gap increases, the motion magnitude and uncertainty grows, leading to an increase in error.
Track optimization. Fig. 15 shows a top-down view of Fig 3. We tried multiple stereo depth methods when developing our system, e.g., BiDAStereo [37] but it still shows jitter and drift (left). That said, as more advanced stereo methods become available, we can adopt them. Ablations (right): w/o $\mathcal{L}_{\mathsf{static}}$ (static content reduces jitter from $\mathcal{L}_{\mathsf{dynamic}}$ but still drifts); w/o $\mathcal{L}_{\mathsf{dynamic}}$ (dynamic trajectories are distorted along camera rays by $\mathcal{L}_{\mathsf{static}}$).
Effect of time gap for DynaDUSt3R. We evaluate DynaDUSt3R’s capability with regard to time gap $\Delta_{t}=t_{1}-t_{0}$ of input images. Fig. 14 shows motion error vs. input frame gap, across 1k Stereo4D pairs. As the time gap increases, the motion magnitude and uncertainty grows, leading to an increase in error.
Figure 15: Track optimization ablation, extending Fig. 3. SOTA video stereo method, BiDAStereo [37], still shows jitter and drift (left). Ablations (right): w/o $\mathcal{L}_{\mathsf{static}}$ static content reduces jitter from $\mathcal{L}_{\mathsf{dynamic}}$ but still drifts; w/o $\mathcal{L}_{\mathsf{dynamic}}$ dynamic trajectories are distorted along camera rays by $\mathcal{L}_{\mathsf{static}}$.
DynaDUSt3R motion head ablation. DynaDUSt3R uses a separate motion head to predict motion for the pointmaps $M^{v\rightarrow t_{q}}$. Alternatively, one can also predict the deformed points with the same point head by conditioning on the time embedding. We compare 1) ours: predicting $M^{v\rightarrow t_{q}}$ with motion heads; and 2) directly regressing $P^{v\rightarrow t_{q}}$ with point heads and evaluate on the same Stereo4d test set in Table 1. Using a single head to predict motion (2) results in a drop in motion accuracy across all metrics: $\text{EPE}_{\text{3D}}$$\downarrow$ = $(\textbf{0.1110}\rightarrow 0.1401)$, $\delta_{\text{3D}}^{0.05}$$\uparrow$ = $(\textbf{65.07}\rightarrow 59.19)$, $\delta_{\text{3D}}^{0.10}$$\uparrow$ = $(\textbf{75.18}\rightarrow 69.73)$. This is likely due to decreased decoder capacity.
@@ 9.1 More results on held-out Stereo4D examples
Fig. 16 shows additional DynaDUSt3R predictions on the Stereo4D held-out test set, extending Fig. 7 from the main paper. Fig. 17 shows additional qualitative examples of motion comparisons on Stereo4D test set, extending Fig. 8 from the main paper. Fig. 17 compares variants of Dyna-DUSt3R trained on different data sources. The model trained on PointOdyssey incorrectly predicts large 3D motions, while the model trained on Stereo4D makes more accurate motion predictions, closer to ground truth.
9.2 Additional track optimization examples